A 50-kg person stands 1.5 m away from one end of a uniform 6.0-m-long scaffold of mass 70.0 kg. Find the tensions in the two vertical ropes supporting the scaffold.


1
Expert's answer
2021-12-20T10:29:36-0500

for sum of torques:


"\sum{\tau_a=0}:1.5\cdot50\cdot9.8+3\cdot70\cdot9.8-6T_2=0\to T_2=465.5\ (N)"


"\sum{\tau_b=0}:3\cdot70\cdot9.8+4.5\cdot50\cdot9.8-6T_1=0\to T_1=710.5\ (N)"